# Exploring Scene Coherence for Semi-Supervised 3D Semantic Segmentation

### v1

@@ Introduction
where $p_{ij}^{l}$ denotes the $j$-th point patch from the sequence. From all labeled scenes, we then construct the patch pool $\mathcal{B}_{\text{Pat}}^{l}$, which can be defined as:
@@
Based on the patch pool, we can obtain the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ by assigning different class labels to each patch, which contains instance information for different categories (e.g., car and person). Following the same steps used for labeled scenes, we further create a pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$ and a pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ from the erased scene $\hat{x}_{i}^{u}$. Unlike $\mathcal{B}_{\text{Pat}}^{l}$ and $\mathcal{B}_{\text{Ins}}^{l}$ covering the whole training set $D_{l}$, the pseudo pools $\mathcal{B}_{\text{Pat}}^{u}$ and $\mathcal{B}_{\text{Ins}}^{u}$ only contain scene and instance information from the current batch and are reset in each iteration. This prevents low-quality pseudo labels from persisting in pools and misleading model training. Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool, ensuring effective inter-scene information transfer.
@@
MixPatch. To model the inter-scene correlation and then guide the transfer of annotation information between labeled scenes and unlabeled scenes, we combine multiple scenes using patch pools, as shown in Figure 4. Figure 4 (a) illustrates the patch mixing (MixPatch), while Figure 4 (b) presents the instance filling (InsFill), which is introduced in the next paragraph. During a training iteration, the mini-batch is comprised of labeled and unlabeled scenes. After the unlabeled scene $x_{i}^{u}$ is transformed into the erased scene $\hat{x}_{i}^{u}$ using the point erasure strategy, we integrate information from the patch pool into $\hat{x}_{i}^{u}$. The patch mixing process can be described as follows:
@@
where $\mathcal{M}$ is the same as defined in Eq. (5), thereby ensuring the full utilization of scene information from the unlabeled scenes. The binary mask $\mathcal{M}$ and its complement $1-\mathcal{M}$ are used to combine inter-scene point cloud patches with the erased scene, facilitating the transfer of annotation information between labeled and unlabeled scenes. To alleviate bias in the patch mixing process, we apply a uniform sampling strategy on the patches of each scene, ensuring that each patch has an equal opportunity to be selected in the mixed scene.
@@
InsFill. In addition to patch mixing, we utilize the instances stored in the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ to further enhance the scene $\hat{x}_{i}^{up}$. Given a patch with location index $j$ in the scene, we randomly select an instance from the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ that was extracted from the $j$-th patch. Patches with identical indices typically correspond to similar LiDAR spatial positions, ensuring consistency between the selected instance and its fill positions in the scene. As depicted in Figure 4 (b), if the selected instance overlaps with existing instances in the scene or lacks sufficient contextual information, the filling operation is bypassed. This judgment process can be easily implemented using the maximum 3D bounding boxes. This data augmentation provides single-instance semantic information for the point cloud scene, enabling the augmented scene $\hat{x}_{i}^{upf}$ and corresponding labels $\hat{y}_{i}^{upf}$ to effectively support semi-supervised training. For the labeled branch, we also apply the same filling strategy as described, selecting instances from the pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ to fill the current scene. This process yields the scene $x_{i}^{lpf}$ along with its corresponding label $y_{i}^{lpf}$.
Model Training
@@
Following the standard teacher-student mutual learning paradigm (Kong et al. 2023b; Sohn et al. 2020), we construct a mini-batch of data with an equal composition of labeled and unlabeled scenes at each iteration. The unlabeled scenes are assigned with pseudo-labels by the teacher model and then train the student model jointly with labeled scenes. We use three types of loss functions to supervise the model training here, which is shown as follows:
$\ \mathcal{L}_{s}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(x_{i}^{l}),y_{i}^{l}),$ (7)
$\ \mathcal{L}_{l}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(x_{i}^{lpf}),y_{i}^{lpf}),$ (8)
$\ \mathcal{L}_{u}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(\hat{x}_{i}^{upf}),\hat{y}_{i}^{upf}),$ (9)
@@
where $\mathcal{L}_{s}$, $\mathcal{L}_{l}$, and $\mathcal{L}_{u}$ denote the segmentation loss for the labeled scene $x_{i}^{l}$, augmented labeled scene $x_{i}^{lpf}$, and augmented unlabeled scene $\hat{x}_{i}^{upf}$, respectively, while $\mathcal{L}_{seg}$ is the original loss of the segmentation network $\Phi(\cdot)$. We sum up all the losses together as our final training target that
@@
where $\lambda_{u}$ and $\lambda_{l}$ are loss weights that balance the two loss terms. After that, we update the teacher model weights via the EMA strategy described in Eq. (1).
SemanticKITTI nuScenes
Method Venue 1% 5% 10% 20% 40% 50% 1% 10% 20% 50%
Cylinder3D CVPR 2021 45.4 45.4 56.1 57.8 58.7 58.7 50.9 65.9 66.6 71.2
MeanTeacher NeurIPS 2017 45.4 - 57.1 59.2 - 60.0 51.6 66.0 67.1 71.7
CBST ECCV 2018 48.8 - 58.3 59.4 - 59.7 53.0 66.5 69.6 71.6
CPS CVPR 2021 46.7 - 58.7 59.6 - 60.5 52.9 66.3 70.0 72.5
GPC ICCV 2021 - 41.8 49.9 58.8 59.9 - - - - -
ScribbleKITTI CVPR 2022 - 49.9 58.7 59.1 60.9 - - - - -
LiM3D CVPR 2023 58.4 59.5 62.2 63.1 63.3 63.6 - - - -
CylinderMix* AAAI 2024 51.3 56.5 60.8 62.0 62.7 62.5 55.7 69.5 72.0 73.3
LaserMix CVPR 2023 50.6 56.7 60.0 61.9 62.3 62.3 55.3 69.9 71.8 73.2
Ours (CoScene) - 54.5 62.1 63.3 63.7 63.8 64.3 56.6 70.2 72.8 73.9
Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods. The results are reported with mIoU scores (%) and the best score for different split is marked in bold. “*” denotes that we reproduce the CylinderMix in the LiDAR branch (Chen et al. 2024) for fair comparison.
Experiments
Datasets and Evaluation Metrics
@@
SemanticKITTI. Following state-of-the-art methods (Kong et al. 2023b; Chen et al. 2024), we evaluate our CoScene on the SemanticKITTI dataset (Behley et al. 2019), which is a large-scale segmentation dataset consisting of LiDAR data derived from the popular KITTI dataset (Geiger, Lenz, and Urtasun 2012). It is composed of 22 sequences, with sequences 00-10 serving as the train set (of which 08 as the val set) and sequences 11-21 as the test set. Moreover, all the point cloud scenes are collected from real-world rural and urban road environments in Germany, and the semantic class labels are mapped to 19 classes in our experiments.
@@
nuScenes. The nuScenes dataset (Fong et al. 2022) is another well-known large autonomous driving dataset. This dataset is composed of 1000 sequences of 20s duration (850 for training and validation, 150 for testing), which are captured by the LiDAR sensor from Boston and Singapore. Following the official splitting protocol, the train and val sets contains 28,130 and 6,019 point cloud scenes, respectively, including per-points annotation for 16 semantic classes.
@@
Evaluation Metrics. For both the SemanticKITTI and nuScenes datasets, we sample different ratios of annotated train scenes to obtain semi-supervised data and ensure the sampling indices are the same as in LaserMix(Kong et al. 2023b) for fair comparison. Meanwhile, we report the performance on the validation set using the mean Intersection over Union (mIoU) metric as the previous works (Li, Shum, and Breckon 2023; Zhu et al. 2021).
Implementation Details
@@
We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D (Contributors 2020). Inspired by the GT-sampling strategy (Yan, Mao, and Li 2018), we set $\tau_{min}$ to 5 in our experiments. For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we use the same parameter setting as in LaserMix. To facilitate patch-based data augmentation, we restrict our focus to point clouds within the X and Y dimensions in the range of $[-50,50]m$ for both the SemanticKITTI and nuScenes datasets. Similar to mainstream work, we also adopt the consistency loss proposed in MeanTeacher (Tarvainen and Valpola 2017) to align the predictions of the teacher and student networks. Additionally, the weights of the consistency loss are assigned as $1\times 10^{3}$ and $2\times 10^{3}$ for the SemanticKITTI and nuScenes datasets, respectively. The weights for $\lambda_{u}$ and $\lambda_{l}$ are set as 1 for both datasets.
Main Results
@@
In this subsection, we compare our method with current SOTA semi-supervised point cloud segmentation methods, using sub-sampling of the datasets under various partition protocols (Kong et al. 2023b). Specifically, in the SemanticKITTI dataset (Behley et al. 2019), we split the entire training set into labeled subsets with ratios of 1%, 5%, 10%, 20%, 40%, and 50%. Similarly, in the nuScenes dataset (Fong et al. 2022), we utilize labeled subsets with ratios of 1%, 10%, 20%, and 50%. We provide detailed comparative results on both datasets against different baseline methods.
Results on SemanticKITTI Dataset. We present the comparison results between our method and state-of-the-art methods on the widely adopted SemanticKITTI dataset in Table 1. It can be clearly seen that our method effectively exploits unlabeled scenes compared to the fully supervised learning baseline Cylinder3D (Zhu et al. 2021), with a dramatic performance boost. Specifically, in the 5% ratio, our approach surpasses the fully supervised baseline by 16.7%. In other settings, our method also shows consistent improvements ranging from 5.6% to 7.2% for ratios of 10%, 20%, and 50%. In addition, comparison to the recently proposed SOTA method LaserMix (Kong et al. 2023b), which is considered a strong SSL-based point cloud segmentation method, our approach achieves improvements of 2% to 6% in all cases. The results also demonstrate that our approach outperforms the current SOTA LiM3D (Li, Shum, and Breckon 2023) in the majority of settings. In some partition protocols, our approach shows superior performance compared to the LiM3D with fewer labeled scenes. For example, our approach trained with 10% labeled scenes outperforms LiM3D trained with 40% labeled scenes. It is worth mentioning that our approach utilizes uniform sampling, similar to LaserMix, to acquire labeled scenes, without utilizing the keyframe sampling method used in LiM3D (Li, Shum, and Breckon 2023), which typically yields higher gains when the amount of labeled data is relatively low.
Results on nuScenes Dataset. As shown in Table 1, our approach still shows a significant improvement in performance compared to state-of-the-art methods on the comprehensive nuScenes dataset (Fong et al. 2022). For instance, our approach outperforms LaserMix (Kong et al. 2023b) by 1.3% and 1.0% at 1% and 20% ratios, respectively, demonstrating the generalization ability of our method.
Ablation Study
In this subsection, we study the effects of intra-scene consistency assurance through the point erasure strategy (PtErase), as well as the exploration of inter-scene correlation, incorporating patch-based mix (MixPatch) and patch-based instance filling data augmentation (InsFill) in our method. Without loss of generalizability, all experiments are conducted on the SemanticKITTI dataset with 1%, 10%, 20% and 50% ratio, and evaluated on val set. Table 2 demonstrates the improvements achieved by each of these components.
Ours CoScene Splits on SemanticKITTI Avg.

### v2

@@ 3.3 Patch-based Data Augmentation
where $p_{ij}^{l}$ denotes the $j$-th point patch from the sequence. For all labeled scenes, by perform the operation in Eq. 3, the patch pool $\mathcal{B}_{\text{Pat}}^{l}$ is constructed, which can be defined as:
@@
Based on the patch pool, we can obtain the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ by assigning different class labels to each patch, as depicted in Fig. 3, which contains instance information for different categories (e.g., car and person).
Following the same steps used for labeled scenes, we further create a pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$ and a pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ from the erased scene $\hat{x}_{i}^{u}$. Unlike $\mathcal{B}_{\text{Pat}}^{l}$ and $\mathcal{B}_{\text{Ins}}^{l}$ covering the whole training set $D_{l}$, the pseudo pools $\mathcal{B}_{\text{Pat}}^{u}$ and $\mathcal{B}_{\text{Ins}}^{u}$ only contain scene and instance information from the current batch and are reset in each iteration. This prevents low-quality pseudo labels from persisting in pools and misleading model training. Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool, ensuring effective semantic diversity.
@@
MixPatch. To model the inter-scene correlation and then guide the transfer of annotation information between labeled scenes and unlabeled scenes, we combine multiple scenes using patch pools. Fig. 4 (a) illustrates the patch mixing (MixPatch) strategy, while Fig. 4 (b) presents the instance filling (InsFill) strategy, which is introduced in the next paragraph. During a training iteration, the mini-batch is comprised of labeled and unlabeled scenes. After the unlabeled scene $x_{i}^{u}$ is transformed into the erased scene $\hat{x}_{i}^{u}$ using the point erasure strategy, we integrate information from the patch pool into $\hat{x}_{i}^{u}$ by MixPatch. The MixPatch process can be described as follows:
@@
where $\mathcal{M}$ is the same as defined in Eq. 5, thereby ensuring the full utilization of scene information from the unlabeled scenes. The binary mask $\mathcal{M}$ and its complement $1-\mathcal{M}$ are used to merge inter-scene point cloud patches with the erased scene, facilitating the generation of diverse semantic information for labeled and unlabeled scenes. To alleviate bias in the patch mixing process, we apply a uniform sampling strategy to the patches in each scene, ensuring that each patch has an equal opportunity of being selected in the mixed scene.
@@
InsFill. In addition to patch mixing, we utilize the instances stored in the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ to further enhance the scene $\hat{x}_{i}^{up}$. Given a patch with location index $j$ in the scene, we randomly select an instance from the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ that was extracted from the $j$-th patch. Patches with identical indices typically correspond to similar LiDAR spatial layouts, ensuring consistency between the selected instance and its fill positions in the scene. As depicted in Fig. 4 (b), if the selected instance overlaps with existing instances in the scene or lacks sufficient contextual information, the filling operation is bypassed. This judgment process can be easily implemented using the maximum 3D bounding boxes. This data augmentation provides instance-level semantic information for the point cloud scene, enabling the augmented scene $\hat{x}_{i}^{upf}$ and corresponding labels $\hat{y}_{i}^{upf}$ to effectively support semi-supervised training. For the labeled branch, we also apply the same filling strategy as described, selecting instances from the pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ to fill the current scene. This InsFill process yields the scene $x_{i}^{lpf}$ along with its corresponding label $y_{i}^{lpf}$.
Method Venue SemanticKITTI [2] Avg. nuScenes [10] Avg.
1% 10% 20% 50% 1% 10% 20% 50%
Meanteacher [38] NeurIPS’2017 45.4 57.1 59.2 60.0 55.4 51.6 66.0 67.1 71.7 64.1
CBST [56] ECCV’2018 48.8 58.3 59.4 59.7 56.6 53.0 66.5 69.6 71.6 65.2
CPS [4] CVPR’2021 46.7 58.7 59.6 60.5 56.4 52.9 66.3 70.0 72.5 65.4
LaserMix (Range View) [15] CVPR’2023 43.4 58.8 59.4 61.4 55.8 49.5 68.2 70.6 73.0 65.3
LaserMix (Voxel) [15] CVPR’2023 50.6 60.0 61.9 62.3 58.7 55.3 69.9 71.8 73.2 67.6
GPC [13] ICCV’2021 54.1 62.0 62.5 62.8 60.4 - - - - -
IGNet [40] WACV’2024 49.0 61.3 63.1 64.8 59.6 - - - - -
LiM3D [23] CVPR’2023 58.4 62.2 63.1 63.6 61.8 - - - - -
LASS3D [22] ECCV’2024 58.5 63.0 64.1 64.5 62.5 - - - - -
DDSemi [21] CVPR’2024 59.3 65.1 66.3 67.0 64.4 58.1 70.2 74.0 76.5 69.7
Ours (AIScene) - 61.2 66.3 67.4 67.9 65.7 63.4 73.7 75.0 76.3 72.1
Improvement $\uparrow$ - +1.9 +1.2 +1.1 +0.9 +1.3 +5.3 +3.5 +1.0 -0.2 +2.4
Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods. The results are reported with mIoU scores (%). The blod and underline scores respectively denote the best and second results.
3.4 Model Training
@@
Following the standard teacher-student mutual learning paradigm [38, 15, 34], we construct a mini-batch of data with an equal composition of labeled and unlabeled scenes at each iteration. The unlabeled scenes are assigned with pseudo-labels by the teacher model and then train the student model jointly with labeled scenes. We use three types of loss functions to supervise the model training here, which is shown as follows:
$\ \mathcal{L}_{s}=\mathcal{L}_{seg}(\Phi_{\theta_{\mathbbm{s}}}(x_{i}^{l}),y_{i}^{l}),$ (7)
$\ \mathcal{L}_{l}=\mathcal{L}_{seg}(\Phi_{\theta_{\mathbbm{s}}}(x_{i}^{lpf}),y_{i}^{lpf}),$ (8)
$\ \mathcal{L}_{u}=\mathcal{L}_{seg}(\Phi_{\theta_{\mathbbm{s}}}(\hat{x}_{i}^{upf}),\hat{y}_{i}^{upf}),$ (9)
@@
where $\mathcal{L}_{s}$, $\mathcal{L}_{l}$, and $\mathcal{L}_{u}$ denote the segmentation loss for the labeled scene $x_{i}^{l}$, augmented labeled scene $x_{i}^{lpf}$, and augmented unlabeled scene $\hat{x}_{i}^{upf}$, respectively, while $\mathcal{L}_{seg}$ is the original loss of the segmentation network $\Phi(\cdot)$. We sum up all the loss term to define the final training objective as:
@@
where $\lambda_{u}$ and $\lambda_{l}$ are loss weights that balance the two loss terms. After that, we update the teacher model weights using the EMA strategy described in Eq. 1.
4 Experiments
4.1 Datasets and Evaluation Metrics
@@
SemanticKITTI. Following state-of-the-art methods [15, 21], we evaluate our AIScene on the SemanticKITTI dataset [2], which is a large-scale segmentation dataset consisting of LiDAR data derived from the popular KITTI dataset [11]. It is composed of 22 sequences, with sequences 00-10 serving as the train set (of which 08 as the val set) and sequences 11-21 as the test set. Moreover, all the point cloud scenes are collected from real-world rural and urban road environments in Germany, and the semantic class labels are mapped to 19 classes in our experiments.
@@
nuScenes. The nuScenes dataset [10] is another well-known large autonomous driving dataset. This dataset is composed of 1000 sequences of 20s duration (850 for training and validation, 150 for testing), which are captured by the LiDAR sensor from Boston and Singapore. Following the official splitting protocol, the train and val sets contain 28,130 and 6,019 point cloud scenes, respectively, including per-points annotation for 16 semantic classes.
@@
Evaluation Metrics. For both the SemanticKITTI and nuScenes datasets, we sample different ratios of annotated train scenes to obtain semi-supervised data. Meanwhile, we report the performance on the validation set using the mean Intersection over Union (mIoU) metric [21, 23].
4.2 Implementation Details
@@
We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D [8]. Inspired by the GT-sampling strategy [47], we set $\tau_{min}$ to 5 in our experiments. For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we use the same parameter setting as in LaserMix. We use the MinkowskiNet [7] and Cylinder3D [55] as the backbone network due to its widely adopted in previous methods [15, 27, 41]. To facilitate patch-based data augmentation, we restrict our focus to point clouds within the X and Y dimensions in the range of $[-50,50]m$ for both the SemanticKITTI and nuScenes datasets. Similar to mainstream work, we also adopt the consistency loss proposed in MeanTeacher [38] to align the predictions of the teacher and student networks. Additionally, the weights of the consistency loss are assigned as $1\times 10^{3}$ and $2\times 10^{3}$ for the SemanticKITTI and nuScenes datasets, respectively. The weights for $\lambda_{u}$ and $\lambda_{l}$ are set as 1 for both datasets. Our method excludes all forms of model ensemble and test-time augmentation during the evaluation.
4.3 Main Results
@@
In this subsection, we compare our method with current SOTA semi-supervised point cloud segmentation methods (e.g., LaserMix [15], LiM3D [23] and DDSemi [21]), using sub-sampling of the datasets under various partition protocols [21]. Specifically, we split the entire training set into labeled subsets with ratios of 1%, 10%, 20%, and 50% in the SemanticKITTI [2] and nuScenes dataset [10]. For further comparison, we also include several classical semi-supervised learning methods [38, 56, 4] from the 2D image domain adapted to 3D segmentation.
Tab. 1 summarizes the quantitative results on SemanticKITTI and nuScenes datasets. Our method consistently outperforms other comparative methods on both datasets, achieving average mIoU improvements of 1.3% and 2.4%, respectively. Notably, the proposed method demonstrates even greater gains on smaller sampling splits, such as 1% and 10%, which are particularly beneficial for the semi-supervised setting. Specifically, with a 1% ratio, our method shows mIoU improvements of up to 1.9% and 5.3% on the two datasets, respectively, compared to the recently proposed SOTA method, DDSemi [21]. These results also validate that our method is less sensitive to reduced labeled data sampling than other methods. Furthermore, we present qualitative results in Fig. 5. As indicated by the purple and red dashed circles, our method achieves better segmentation of boundary regions and small objects.
4.4 Ablation Study
In this subsection, We first validate the adaptability of our method to different segmentation models. We then examine the impact of intra-scene consistency through the point erasure strategy and explore inter-scene correlation by incorporating patch-based mixing and patch-based instance filling data augmentation.
Figure 5: Qualitative comparison results from LiDAR bird’s eye view on SemanticKITTI val set. We set the point with correct and incorrect predicted semantic class in blue and orange for better visualization. Best viewed in color.
Adaptability with Different Segmentation Methods. Our AIScene is segmentor-agnostic and can be easily used to further improve off-the-shelf 3D fully-supervised segmentors under the semi-supervised setting. Tab. 2 shows that the proposed method can substantially boost the performance with Cylinder3D [55] as the backbone, indicating the generalization ability of our framework. For consistency in comparison, the subsequent experiments are based on Cylinder3D unless otherwise specified.
Effectiveness of Point Erasure Strategy. The 1st row in Tab. 3 reports the performance of the supervision-only Cylinder3D model [55], which serves as the baseline. When PtErase is applied to the baseline, it effectively boosts performance by ensuring consistency during forward and backward propagation, thus addressing the limitations of the pseudo-label mechanism. This consistency is achieved through a plug-and-play point erasure strategy, which can be easily integrated into methods relying on the pseudo-label mechanism. To further validate the importance of maintaining consistency during forward and backward propagation, we embedded the PtErase strategy into LaserMix [15]. As illustrated in Fig. 6 (a), using the PtErase strategy improves performance across all ratios, underscoring the significance of intra-scene consistency and demonstrating the generalizability of our method. Moreover, Fig. 6 (b) shows that the proportion of points removed by the erasure strategy is initially high at the start of training, but gradually decreases as the network converges. Furthermore, We present a comparison of class probability distributions in Fig. 7 to validate the effectiveness of the point erasure strategy. The results demonstrate that this strategy reduces classification ambiguity, enabling the model to better capture the features of the correct class.
Datasets Methods 1% 10% 20% 50% Avg.
Cylinder3D [55] 45.4 56.1 57.8 58.7 54.5
Ours (AIScene) 54.5 63.3 63.7 64.3 61.5
SemanticKITTI Improvement +9.1 +7.2 +5.9 +5.6 +7.0
Cylinder3D [55] 50.9 65.9 66.6 71.2 63.7
Ours (AIScene) 56.6 70.2 72.8 73.9 68.4
nuScenes Improvement +5.7 +4.3 +6.2 +2.7 +4.7
Table 2: Ablation study on the adaptability of our AIScene using Cylinder3D as the segmentation model.
Ours AIScene Splits on SemanticKITTI Avg.